OpenStudy (noahbred):

Help me understand please Convert the following measurements. (Remember both the direction and number of decimal “hops” to use; all questions have multiple hops.) 1. 0.25 mm = __________ m 0.025 2.5 0.00025 250

OpenStudy (wcrmelissa2001):

OK so first things first. How many mm is there in m? (Don't act smart and tell me there are 2...)

OpenStudy (wcrmelissa2001):

there is 100 cm in 1 metre. And 10 millimeters in 1 cm Hence in 1 metre there are 10*100=1000mm

OpenStudy (noahbred):

hahahaha umm 7 idk

OpenStudy (wcrmelissa2001):

you understand the conversion I wrote??

OpenStudy (noahbred):

oh didnt see it sorry i sec

OpenStudy (noahbred):

ohhh theres 3

OpenStudy (wcrmelissa2001):

what do you mean there's 3mm in 1 metre?

OpenStudy (noahbred):

idk? i am very confused ?

OpenStudy (wcrmelissa2001):

Hm... Ok so when you convert mm to m. You first need to know how many mm there is in 1 metre. Just like how you need to know how many cm there is in 1m if you are converting metres to centimetres or vice versa. So in this case, there is 1000mm in 1 metre (because there is 10mm in 1 cm. And there is 100 cm in 1 metre. Hence it's 10*100=1000mm in 1 metre) Let's start with a simpler question k? let's say it's 2000mm convert to metre. The steps would be to first recognise that 1000 mm = 1m Therefore 2000mm (which equals 1000*2) = 1*2 m So 2000mm = 2m Apply this concept to the question (I hope you understand)

OpenStudy (noahbred):

Thanks!